how do i find the equations of these roots? 6,(-2)/(3)?
Answer by scott8148(6628)

if r is a root, then x-r is a factor

(x - 6)(x + 2/3) = 0

FOILing ___ x^2 - (16/3)x - 4 = 0

multiplying by 3 to clear fraction ___ 3x^2 - 16x - 12 = 0